# Continual Test-time Adaptation for End-to-end Speech Recognition on Noisy Speech

### v1

@@ Abstract
Deep learning-based end-to-end automatic speech recognition (ASR) has made significant strides but still struggles with performance on out-of-domain (OOD) samples due to domain shifts in real-world scenarios. Test-Time Adaptation (TTA) methods address this issue by adapting models using test samples at inference time. However, current ASR TTA methods have largely focused on non-continual TTA, which limits cross-sample knowledge learning compared to continual TTA. In this work, we propose a Fast-slow TTA framework for ASR, which leverages the advantage of continual and non-continual TTA. Within this framework, we introduce Dynamic SUTA (DSUTA), an entropy-minimization-based continual TTA method for ASR. To enhance DSUTA’s robustness on time-varying data, we propose a dynamic reset strategy that automatically detects domain shifts and resets the model, making it more effective at handling multi-domain data. Our method demonstrates superior performance on various noisy ASR datasets, outperforming both non-continual and continual TTA baselines while maintaining robustness to domain changes without requiring domain boundary information.
@@
Guan-Ting Linthanks: Equal contribution., Wei-Ping Huang, Hung-yi Lee National Taiwan University, Taiwan {f10942104, hungyilee}@ntu.edu.tw, thomas1232121@gmail.com
@@ 1 Introduction
Deep learning-based end-to-end automatic speech recognition (ASR) has made remarkable progress in recent years, achieving low recognition error rates for in-domain samples. However, domain shifts frequently occur in real-world scenarios. Although recent large-scale ASR models exhibit some generalization to out-of-domain (OOD) test instances, their performance on OOD samples still lags behind that of in-domain samples.
To address domain shift issues during inference, Test-Time Adaptation (TTA) is an attractive method. It adapts the model using only single or batched test samples without needing the source training data at testing time. Specifically, TTA adapts the source model through unsupervised objectives like Entropy Minimization (EM)  (Wang et al., 2020) or Pseudo-Labeling (PL) (Goyal et al., 2022) in inference time.
TTA methods initially gained prominence in the computer vision field. Non-continual TTA methods adapt the source model for each test utterance and reset to the original model for subsequent samples (Wang et al., 2020), whereas Continual TTA (CTTA) continuously adapts the model for target domains, leveraging knowledge learned across samples to improve performance (Niu et al., 2022a, b; Press et al., 2024).
Figure 1: Illustration of the proposed Fast-slow TTA framework and dynamic reset strategy with time-varying speech domains. The Fast-slow TTA framework can leverage the advantages of continual and non-continual TTA, and the dynamic reset strategy automatically detects domain shifts and resets the model to the source model during testing.
In speech recognition, recent studies have tailored TTA methods with entropy-minimization-based optimization (Lin et al., 2022; Kim et al., 2023; Liu et al., 2023), proposing various training objectives and demonstrating effectiveness on different datasets. However, existing ASR TTA methods focus on non-continual TTA, limiting the ability to learn knowledge across samples. There is limited research on CTTA for end-to-end ASR. Recently, AWMC (Lee et al., 2023) proposed a pseudo-labeling CTTA method for ASR on single test domain. However, as shown in previous work (Lin et al., 2022), pseudo-labeling is not as effective as EM-based methods, and its ability on long multi-domain testing data is unknown.
@@
We demonstrate the effectiveness of our method on single-domain and multi-domain time-varying ASR benchmarks under different acoustic conditions, simulating real-world changing environments. Our method outperforms the strong single-utterance baseline SUTA Lin et al. (2022) and the CTTA baseline AWMC Lee et al. (2023), showing robustness to domain changes even without knowing the domain boundaries.
@@ 2 Related Works
Non-continual TTA for ASR: non-continual TTA methods adapt the source model for each test utterance and reset to the original model for subsequent samples. SUTA Lin et al. (2022) introduces the first TTA approach for non-autoregressive ASR, based on entropy minimization and minimum class confusion. SGEM Kim et al. (2023) extends TTA to autoregressive ASR models by introducing a general form of entropy minimization.  Liu et al. (2023) enhances TTA with confidence-enhanced entropy minimization and short-term consistency regularization. However, these non-continual TTA methods views each utterance independently, which only relies on a single utterance and fails to leverage the knowledge across a stream of test samples to improve the adaptation.
Continual TTA: Unlike non-continual TTA, which resets to the source model for each sample, continual TTA enables the online model to use learned knowledge to handle gradual changes in the target domain. However, it may suffer from model collapse if adaptation is unstable when the data stream is too long. To improve CTTA’s performance and stability, studies in the computer vision field have developed solutions like stochastic model restoring (Wang et al., 2022), sample-efficiency entropy minimization (Niu et al., 2022a), sharpness-aware reliable entropy minimization Niu et al. (2022b), and fixed frequency model reset (Press et al., 2024).
In the ASR research, there are limited studies on CTTA ASR. Recently, AWMC Lee et al. (2023) attempts continual TTA on ASR using a pseudo-labeling approach with an extra anchor model to prevent model collapse. However, AWMC Lee et al. (2023) only measures the performance on single-domain data with the pseudo-labeling method. This work focuses on multi-domain time-varying long data streams. We propose a fast-slow TTA framework and dynamic reset strategy based on an entropy minimization-based CTTA method, which achieves better performance and stability.
@@ 3.1 Fast-slow TTA Framework
Non-continual TTA treats each sample as an independent learning event. The adaptation process can fit the current sample without affecting future samples; however, the learned knowledge cannot be transferred to future samples. In contrast, continual TTA can utilize the learned knowledge, but fitting the current sample might severely affect TTA performance on future samples. For instance, if the model overly fits the current sample and (model collapse), the performance on future samples will significantly degrade with continual TTA, whereas in non-continual TTA, the performance remains unaffected.
@@
Fast-slow TTA generalizes non-continual TTA. If $U(\phi_{t},x_{t})=\phi_{t}$, i.e., $\phi_{t}$ remains constant over time, the framework degenerates to non-continual TTA. To the best of our knowledge, this is the first time such an approach has been applied to TTA for ASR.
@@ 3.2 Dynamic SUTA
We propose Dynamic SUTA (DSUTA), a fast-slow TTA method based on SUTA (Lin et al., 2022). SUTA is the very first method for ASR TTA. Given pre-trained parameters $\phi_{pre}$, for every incoming sample $x_{t}$, SUTA adapts $\phi_{pre}$ for $N$ steps with the objective $\mathcal{L}_{suta}$. $\mathcal{L}_{suta}$ consists of entropy loss and minimum class confusion loss. Entropy minimization aims to sharpen class distribution, and minimum class confusion aims to reduce the correlation between different prediction classes. See Appendix A.5 for the detailed loss function. Model parameters are reset to $\phi_{pre}$ when the next sample arrives.
For DSUTA, the adaptation algorithm $A$ is set exactly the same as SUTA, which iteratively adapts $\phi_{t}$ for $N$ steps with $\mathcal{L}_{suta}$ on $x_{t}$. To construct the update algorithm $U$, we introduce a small buffer $\mathcal{B}$ with size $M$. For every $M$ steps, the buffer is filled and we calculate $\mathcal{L}_{suta}$ from these $M$ samples to update the meta-parameters $\phi_{t}$ with gradient descent. The buffer is then cleared. Thus, the meta-parameters $\phi_{t}$ gradually evolve by mini-batch gradient descent with batch size $M$. DSUTA can be viewed as a variant of SUTA, which starts the adaptation from dynamically changing $\phi_{t}$ instead of the fixed $\phi_{pre}$. Denote $\mathcal{L}_{suta}(\phi,x)$ as the loss of sample $x$ on model $\phi$. Algorithm 1 describes the pseudo code of DSUTA.
@@ 3.3 DSUTA with Dynamic Reset Strategy
As time progresses and the testing domain changes, multiple domain shifts significantly challenge the robustness of continual TTA methods. Recently, Press et al. (2024) has shown that model reset at a fixed frequency, which resets the current parameters to the pre-trained ones at regular intervals, is a simple yet effective strategy. Therefore, we attempt to incorporate model reset strategy to update the meta-parameters $\phi_{t}$ in DSUTA [^1] Non-continual TTA can be viewed as the case where we apply model reset at every time step.. However, determining the optimal reset frequency in reality is challenging. To automatically determine when to apply model reset to $\phi_{t}$, we propose a dynamic reset strategy, which actively detects large distribution shifts and dynamically resets $\phi_{t+1}=\phi_{pre}$.
@@
3.3.1 Loss Improvement Index
We aim to find an indicator that measures the extent of the distribution shift from the base domain $\mathcal{D}$. To identify an appropriate indicator, we observed that given a model, we observed that given a model $\phi_{\mathcal{D}}$ trained on domain $\mathcal{D}$, $\mathcal{L}_{suta}$ for in-domain samples is empirically lower than that for out-of-domain samples. This suggests that $\mathcal{L}_{suta}(\phi_{\mathcal{D}},x_{t})$ might be a good indicator. Additionally, we found that subtracting the loss from the pre-trained model, $\mathcal{L}_{suta}(\phi_{pre},x_{t})$, is beneficial [^2] See Section 5.1 for more discussion on indicator choice.. This subtraction acts as a normalization to remove the inherent difficulty introduced by the data sample itself. Overall, we define Loss Improvement Index (LII) as our indicator:
@@ 3.3.2 Domain Construction Stage and Shift Detection Stage
Our goal is to estimate the distribution of LII. We construct $\phi_{\mathcal{D}}=\phi_{r+k}$ as the meta-parameters after observing $k$ samples since the last reset. Calculating the LII requires $\phi_{\mathcal{D}}$, and since TTA is an online process, $K-k$ is the number of LIIs we can collect for statistical estimation. A smaller $k$ might not suffice for $\phi_{\mathcal{D}}$ to adequately represent the domain, while a larger $k$ reduces the number of data points we can gather for estimation. Therefore, we empirically set $k=\lfloor\frac{K}{2}\rfloor$.
@@
Using the LII of a single sample for the hypothesis test can be too sensitive. The averaged LII from multiple samples reduces variance and yields more reliable results. With DSUTA, we perform the hypothesis test every $M$ steps, using the $M$ samples in DSUTA’s buffer to calculate the averaged LII. The final shift detection stage is defined as follows:
@@
Here, $\mathcal{B}$ represents the buffer containing the most recent $M$ samples. In our implementation, we further introduce a patience parameter $P$ to enhance the algorithm’s stability. Please refer to Appendix A.6 for the detailed algorithm.
@@ 4.1 Dataset
4.1.1 Single-domain Data
Librispeech-C: we follow previous works (Kim et al., 2023) to add background noises from MS-SNSD (Reddy et al., 2019) into Librispeech test set (Panayotov et al., 2015). The noises include air conditioner (AC), airport announcement (AA), babble (BA), copy machine (CM), munching (MU), neighbors (NB), shutting door (SD), typing (TP), and vacuum cleaner (VC). We also apply Gaussian noise (GS) as in  (Lin et al., 2022), resulting in 10 different noises in total. The Signal-to-Noise Ratio (SNR) is set to 5 dB.333(Kim et al., 2023) reported using 10dB noise but their source code and results show that they use 5dB.
CHiME-3 (Barker et al., 2017): a noisy version of WSJ corpus with real-world environmental noises at 16kHz.
To reduce GPU memory usage, we exclude samples with raw lengths longer than 20 seconds in all experiments. This removes about 1% of the data.
@@ 4.1.2 Multi-domain Time-varying Data
We create three time-changing multi-domain test sequences by concatenating different corruptions from LibriSpeech-C.
@@ 4.2 Baselines
Non-continual TTA Baselines: (a) SUTA (Lin et al., 2022) leverages unsupervised objectives (entropy minimization and minimum class confusion) to reduce uncertainty and minimize class correlations. Temperature smoothing is applied to flatten the output probability distributions, addressing issues with over-confident predictions. The adaptation process involves iteratively minimizing a combined loss of EM and MCC. (b) SGEM (Kim et al., 2023) propose a general form of entropy minimization with negative sampling.
Continual TTA Baselines: (c) CSUTA is a straightforward continual version of SUTA without resetting parameters. (d) AWMC Lee et al. (2023) utilizes the anchor model to generate initial pseudo labels, the chaser model updates itself using these pseudo labels for self-training, and the leader model refines predictions through an exponential moving average.
@@ 4.3 Implementation Details
We use the wav2vec 2.0-base model fine-tuned on Librispeech 960 hours [^4] https://huggingface.co/facebook/wav2vec2-base-960h as the source ASR model. For SUTA, we follow the official implementation [^5] https://github.com/DanielLin94144/Test-time-adaptation-ASR-SUTA, where an additional reweighting trick is applied on the minimum class confusion loss. The default adaptation step of SUTA is $N=10$, as specified in the original paper. For SGEM, we follow the official implementation [^6] https://github.com/drumpt/SGEM. For CSUTA, we set the adaptation step to $N=1$ since we found that any higher value would cause severe model collapse. We re-implemented AWMC with wav2vec 2.0, as there is no official code, and all hyperparameters follow the original paper. For the proposed DSUTA, the default buffer size is $M=5$, and the adaptation step is $N=10$.

### v2

@@ Abstract
Deep Learning-based end-to-end Automatic Speech Recognition (ASR) has made significant strides but still struggles with performance on out-of-domain samples due to domain shifts in real-world scenarios. Test-Time Adaptation (TTA) methods address this issue by adapting models using test samples at inference time. However, current ASR TTA methods have largely focused on non-continual TTA, which limits cross-sample knowledge learning compared to continual TTA. In this work, we first propose a Fast-slow TTA framework for ASR that leverages the advantage of continual and non-continual TTA. Following this framework, we introduce Dynamic SUTA (DSUTA), an entropy-minimization-based continual TTA method for ASR. To enhance DSUTA robustness for time-varying multi-domain data, we design a dynamic reset strategy to automatically detect domain shifts and reset the model. Our method demonstrates superior performance on various noisy ASR datasets, outperforming both non-continual and continual TTA baselines while maintaining robustness to domain changes without requiring domain boundary information [^1] The source code is available at https://github.com/
hhhaaahhhaa/Dynamic-SUTA
@@
Guan-Ting Lin, Wei-Ping Huang, Hung-yi Lee Graduate Institute of Communication Engineering, National Taiwan University Taiwan {f10942104, hungyilee}@ntu.edu.tw, thomas1232121@gmail.com
@@ 1 Introduction
Deep learning-based end-to-end Automatic Speech Recognition (ASR) has made remarkable progress in recent years, achieving low recognition error rates for in-domain samples. However, domain shifts frequently occur in real-world scenarios. Although recent large-scale ASR models exhibit some generalization to out-of-domain test samples, their performance on out-of-domain samples still lags behind the in-domain performance.
Test-time adaptation (TTA) is an attractive method to address domain shift issues during inference time. TTA adapts the model using only single or batched test samples without needing the source training data at testing time. Specifically, the source model is adapted via unsupervised objectives like Entropy Minimization (EM)  (Wang et al., 2020) or Pseudo-Labeling (PL) (Goyal et al., 2022) in inference time. TTA methods can be characterized into two categories: 1) Non-continual TTA methods adapt the source model for each test utterance and reset to the original model for subsequent samples (Wang et al., 2020), and 2) Continual TTA (CTTA) continuously adapts the model for target domains, leveraging knowledge learned across samples to improve performance (Niu et al., 2022a, b; Press et al., 2024).
Figure 1: Illustration of the proposed Fast-slow TTA framework and dynamic reset strategy with time-varying speech domains. The Fast-Slow TTA framework includes meta-parameters that update slowly to capture cross-domain knowledge, while other parameters update fast for the incoming test samples. The Dynamic reset strategy automatically detects domain shifts and resets the model to the source model.
TTA methods initially strive in the field of computer vision (Wang et al., 2020; Niu et al., 2022a, b; Press et al., 2024). In speech recognition, recent studies have tailored TTA methods with EM-based optimization (Lin et al., 2022; Kim et al., 2023; Liu et al., 2023), proposing new training objectives and demonstrating effectiveness across datasets. However, existing ASR TTA methods only focus on non-continual TTA, constraining the model to learn knowledge across samples. There is limited research on CTTA for end-to-end ASR. Recently, AWMC (Lee et al., 2023) proposed a pseudo-labeling CTTA method for ASR on a single test domain. However, as shown in previous work (Lin et al., 2022), pseudo-labeling is not as effective as EM-based methods, and its ability on long multi-domain testing data is unknown.
@@
We demonstrate the effectiveness of our method on single-domain and multi-domain time-varying ASR benchmarks under different acoustic conditions, simulating real-world changing environments. Our method outperforms the strong single-utterance baseline SUTA (Lin et al., 2022) and the CTTA baseline AWMC (Lee et al., 2023), showing robustness to domain changes even without knowing the domain boundaries.
@@ 2 Related Works
2.1 Non-continual TTA for ASR
Non-continual TTA methods adapt the source model for each test utterance and reset to the original model for subsequent samples. SUTA (Lin et al., 2022) introduces the first TTA approach for non-autoregressive ASR, based on entropy minimization and minimum class confusion. SGEM (Kim et al., 2023) extends TTA to autoregressive ASR models by introducing a general form of entropy minimization.  Liu et al. (2023) enhances TTA with confidence-enhanced entropy minimization and short-term consistency regularization. However, these non-continual TTA methods view each utterance independently, which only relies on a single utterance and fails to leverage the knowledge across a stream of test samples to improve the adaptation.
2.2 Continual TTA
Unlike non-continual TTA, which resets to the source model for each sample, continual TTA enables the online model to use learned knowledge to handle gradual changes in the target domain. However, it may suffer from model collapse if adaptation is unstable when the data stream is too long. To improve the performance and stability of CTTA, studies in the computer vision field have developed solutions like stochastic model restoring (Wang et al., 2022), sample-efficiency entropy minimization (Niu et al., 2022a), sharpness-aware reliable entropy minimization (Niu et al., 2022b), and fixed frequency model reset (Press et al., 2024).
In the ASR research, there are limited studies on CTTA ASR. Recently, AWMC (Lee et al., 2023) attempts continual TTA on ASR using a pseudo-labeling approach with an extra anchor model to prevent model collapse. However, AWMC (Lee et al., 2023) only measures the performance on single-domain data with the pseudo-labeling method. This work focuses on multi-domain time-varying long data streams. We propose a fast-slow TTA framework and dynamic reset strategy based on an entropy minimization-based CTTA method, which achieves better performance and stability.
@@ 3.1 Fast-slow TTA Framework
Non-continual TTA treats each sample as an independent learning event. The adaptation process can fit the current sample without affecting future samples; however, the learned knowledge cannot be transferred to future samples. In contrast, continual TTA utilizes learned knowledge, but overfitting the current sample can adversarially degrade performance on future samples. For instance, if the model overly fits the current sample and (model collapse), the performance on future samples will significantly degrade with continual TTA, whereas in non-continual TTA, the performance remains unaffected.
@@
Fast-slow TTA generalizes to continual and non-continual TTA. If $U(\phi_{t},x_{t})=\phi_{t}$, i.e., $\phi_{t}$ remains constant over time, the framework degenerates to non-continual TTA. On the other hand, if $A=U$, i.e. $\phi_{t+1}=\widehat{\phi}_{t}$, the framework degenerates to continual TTA.
@@ 3.2 Dynamic SUTA
We propose Dynamic SUTA (DSUTA), a fast-slow TTA method based on SUTA (Lin et al., 2022). Specifically, given pre-trained parameters $\phi_{pre}$, for every incoming sample $x_{t}$, SUTA adapts $\phi_{pre}$ for $N$ steps with the objective $\mathcal{L}_{suta}$. $\mathcal{L}_{suta}$ consists of entropy loss and minimum class confusion loss. Entropy minimization aims to sharpen class distribution, and minimum class confusion aims to reduce the correlation between different prediction classes. See Appendix A.4 for the detailed loss function. Model parameters are reset to $\phi_{pre}$ when the next sample arrives.
For DSUTA, the adaptation algorithm $A$ is set exactly the same as SUTA, which iteratively adapts $\phi_{t}$ for $N$ steps with $\mathcal{L}_{suta}$ on $x_{t}$. To construct the update algorithm $U$, we introduce a small buffer $\mathcal{B}$ with size $M$. For every $M$ step, the buffer is filled and we calculate $\mathcal{L}_{suta}$ from these $M$ samples to update the meta-parameters $\phi_{t}$ with gradient descent. The buffer is then cleared. Thus, the meta-parameters $\phi_{t}$ gradually evolve by mini-batch gradient descent with batch size $M$. DSUTA can be viewed as a variant of SUTA, which starts the adaptation from dynamically changing $\phi_{t}$ instead of the fixed $\phi_{pre}$. Denote $\mathcal{L}_{suta}(\phi,x)$ as the loss of sample $x$ on model $\phi$. Algorithm 1 describes the pseudo code of DSUTA.
@@ 3.3 DSUTA with Dynamic Reset Strategy
As time progresses and the testing domain changes, multiple domain shifts significantly challenge the robustness of continual TTA methods. Recently, Press et al. (2024) has shown that model reset at a fixed frequency, which resets the current parameters to the pre-trained ones at regular intervals, is a simple yet effective strategy. Therefore, we attempt to utilize model reset strategy to update the meta-parameters $\phi_{t}$ in DSUTA [^2] Non-continual TTA can be viewed as the case where we apply model reset at every time step.. However, determining the optimal reset frequency in reality is challenging. To automatically determine when to apply model reset to $\phi_{t}$, we propose a dynamic reset strategy that actively detects large distribution shifts and dynamically resets $\phi_{t+1}=\phi_{pre}$.
@@
3.3.1 Loss Improvement Index (LII)
We aim to find an indicator that measures the extent of the distribution shift from the base domain $\mathcal{D}$. To identify an appropriate indicator, we observed that given a model $\phi_{\mathcal{D}}$ trained on domain $\mathcal{D}$, $\mathcal{L}_{suta}$ for in-domain samples is empirically lower than that for out-of-domain samples. This suggests that $\mathcal{L}_{suta}(\phi_{\mathcal{D}},x_{t})$ might be a good indicator. Additionally, we found that subtracting the loss from the pre-trained model, $\mathcal{L}_{suta}(\phi_{pre},x_{t})$, is beneficial to normalize the inherent difficulty introduced by the data sample itself [^3] See Section 5.1 for more discussion on indicator choice.. Overall, we define Loss Improvement Index (LII) as our indicator:
@@ 3.3.2 Domain Construction Stage and Shift Detection Stage
The goal is to estimate the distribution of LII. We construct $\phi_{\mathcal{D}}=\phi_{r+k}$ as the meta-parameters after observing $k$ samples since the last reset. Calculating the LII requires $\phi_{\mathcal{D}}$, and since TTA is an online process, $K-k$ is the number of LIIs we can collect for statistical estimation. A smaller $k$ might not suffice for $\phi_{\mathcal{D}}$ to adequately represent the domain, while a larger $k$ reduces the number of data points we can gather for estimation. Therefore, we empirically set $k=\lfloor\frac{K}{2}\rfloor$.
@@
Additionally, using the LII of a single sample for the hypothesis test is too sensitive. The averaged LII from multiple samples reduces variance and yields more reliable results. With DSUTA, we perform the hypothesis test every $M$ step, using the $M$ samples in DSUTA’s buffer to calculate the averaged LII. The final shift detection stage is defined as follows:
@@
Here, $\mathcal{B}$ represents the buffer containing the most recent $M$ samples. In our implementation, we further introduce a patience parameter $P$ to enhance the stability. Please refer to the Appendix Algorithm 2 for details.
@@ 4.1 Dataset
4.1.1 Single-domain Simulated Noisy Data
Corrupted Librispeech (LS-C): we follow previous works (Kim et al., 2023) by adding background noises from MS-SNSD (Reddy et al., 2019) into Librispeech test set (Panayotov et al., 2015). The noises include air conditioner (AC), airport announcement (AA), babble (BA), copy machine (CM), munching (MU), neighbors (NB), shutting door (SD), typing (TP), and vacuum cleaner (VC). We also apply Gaussian noise (GS) as in  (Lin et al., 2022), resulting in 10 different noises in total. The Signal-to-Noise Ratio (SNR) is set to 5 dB.444(Kim et al., 2023) reported using 10dB noise but their source code and results show that they use 5 dB.
@@ 4.1.2 Multi-domain Time-varying Data
We create three time-changing multi-domain test data streams by concatenating different corruptions from LS-C.
@@
4.1.3 Multi-domain Real Noisy Data
CHiME-3 (Barker et al., 2017): a noisy version of WSJ corpus mixed with real speech recorded in four noisy environments (Cafe, Bus, Street, Pedestrian Area). In this work, different types of noisy speech are randomly distributed in a sequence across time.
@@ 4.2 Baselines
4.2.1 Non-continual TTA Baselines
1) SUTA (Lin et al., 2022) leverages unsupervised objectives (entropy minimization and minimum class confusion) to reduce uncertainty and minimize class correlations. Temperature smoothing is applied to flatten the output probability distributions, addressing issues with over-confident predictions. The adaptation process involves iteratively optimizing the objective of entropy minimization and minimal class correlation. 2) SGEM (Kim et al., 2023) propose a general form of entropy minimization with negative sampling.
4.2.2 Continual TTA Baselines
3) CSUTA is a straightforward continual version of SUTA without resetting parameters. 4) AWMC (Lee et al., 2023) utilizes the anchor model to generate initial pseudo labels, the chaser model updates itself using these pseudo labels for self-training, and the leader model refines predictions through an exponential moving average.
@@ 4.3 Implementation Details
We use the wav2vec 2.0-base model fine-tuned on Librispeech 960 hours [^5] https://huggingface.co/facebook/wav2vec2-base-960h as the source ASR model. For SUTA, we follow the official implementation [^6] https://github.com/DanielLin94144/Test-time-adaptation-ASR-SUTA, where an additional reweighting trick is applied on the minimum class confusion loss. The default adaptation step of SUTA is $N=10$, as specified in the original paper. For SGEM, we follow the official implementation [^7] https://github.com/drumpt/SGEM. For CSUTA, we set the adaptation step to $N=1$ since we found that any higher value would cause severe model collapse. We re-implemented AWMC with wav2vec 2.0, as there is no official code, and all hyperparameters follow the original paper. For the proposed DSUTA, the default buffer size is $M=5$, and the adaptation step is $N=10$. To reduce GPU memory usage, we exclude samples with raw lengths longer than 20 seconds in all experiments. This removes about 1% of the data.
For hyperparameter search, we investigate batch sizes (M=3, 5, 10) and domain construction steps (K=50, 100, 200), and find out that our method is robust across different setups. For more details, please see the Appendix A.2 section.